OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + y = 6 6x + 2y = 8

OpenStudy (anonymous):

are you sure you wrote the equations correctly?

OpenStudy (anonymous):

(2)3x+y=6--> 6x+2y=12 (1)6x+2y=8-->6x+2y+8 Same sign you subtract ........ Yes i think you put the question wrong

OpenStudy (anonymous):

or inversely.. y=6-3x 6x + 2(6-3x) = 8 6x + 12 - 6x = 8 12 = 8? Impossible

OpenStudy (anonymous):

Or 4 But <---- Not right Soo I d k